OpenStudy (anonymous):

The line passing through points (x, 4) and (4, -5) is perpendicular to a line with a slope of -7/3. What is the value of x?

OpenStudy (amistre64):

flip it and change it: 3/7 x,4 -4,5 ----- x-4,9 ; 9/(x-4) = 3/7 x-4 = 7(9)/3 x = 21 +4 = 25

OpenStudy (amistre64):

25,4 -4,5 ----- 21,9 ; slope = 9/21 = 3/7 3/7.-7/3 = -1 = good x = 25

OpenStudy (anonymous):

thanks LOVE (: your the best!